OpenStudy (anonymous):

Calculate the integral: Int(sin(10x)*Cos(10x+1)^(1/8),x)

OpenStudy (anonymous):

\[\[\int\limits\limits \sin(10x)\cos(10x+1)^1/8)\] its to the 1/8th power

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=integrate+sin%2810x%29*Cos%2810x%2B1%29^%281%2F8%29 not sure if you do it.

OpenStudy (anonymous):

(Cos(10x+1))^(1/8)

OpenStudy (anonymous):

I don't understand the hypergeometric function. Is this really necessary for a Calc I class? My professor never lectured on it.

OpenStudy (turingtest):

if this is calc I that would strongly suggest there is a typo somewhere

OpenStudy (lalaly):

only the cos is raised to the power 1/8?

OpenStudy (anonymous):

yes